# Thinking in Granularity: Dynamic Quantization for Image Super-Resolution by Intriguing Multi-Granularity Clues

### v1

@@ Abstract
Dynamic quantization has attracted rising attention in image super-resolution (SR) as it expands the potential of heavy SR models onto mobile devices while preserving competitive performance. Existing methods explore layer-to-bit configuration upon varying local regions, adaptively allocating the bit to each layer and patch. Despite the benefits, they still fall short in the trade-off of SR accuracy and quantization efficiency. Apart from this, adapting the quantization level for each layer individually can disturb the original inter-layer relationships, thus diminishing the representation capability of quantized models. In this work, we propose Granular-DQ, which capitalizes on the intrinsic characteristics of images while dispensing with the previous consideration for layer sensitivity in quantization. Granular-DQ conducts a multi-granularity analysis of local patches with further exploration of their information densities, achieving a distinctive patch-wise and layer-invariant dynamic quantization paradigm. Specifically, Granular-DQ initiates by developing a granularity-bit controller (GBC) to apprehend the coarse-to-fine granular representations of different patches, matching their proportional contribution to the entire image to determine the proper bit-width allocation. On this premise, we investigate the relation between bit-width and information density, devising an entropy-to-bit (E2B) mechanism that enables further fine-grained dynamic bit adaption of high-bit patches. Extensive experiments validate the superiority and generalization ability of Granular-DQ over recent state-of-the-art methods on various SR models. Code will be available at https://github.com/MmmingS/Granular-DQ.git.
@@
Figure 1: Visual comparison of (a) previous dynamic quantization pipeline [15] that adapt the bit allocation for layers and patches simultaneously and (b) our Granular-DQ pipeline conducts patch-wise and layer-invariant dynamic quantization, which contains two steps: 1) granularity-aware bit allocation and 2) fine-grained bit-width adaption based on the entropy statistics. Our method recovers a better SR image with a lower average bit.
1 Introduction
@@
Single image super-resolution (SISR) has been a fundamental task in the computer vision community, aiming to recover high-resolution (HR) images from corrupted low-resolution (LR) input. Recently, from the pioneering deep learning-based method [9], convolutional neural networks (CNN) [9, 21, 38, 51, 2, 27] and transformers [30, 33, 50, 6] have dominated SISR. While the SR performance continues to achieve breakthroughs, the model complexity of later methods also increases constantly, which limits their practical applications, especially tackling large-size images (e.g. 2K and 4K). This raises interest in compressing deep SR models to unlock their potential on resource-constrained devices.
@@
Model quantization [55, 7] has emerged as a promising technology that reduces both computational overhead and memory cost with minimal performance sacrifice, where the effectiveness has been demonstrated in a wide range of high-level tasks [55, 7, 3, 5]. Some prior works design SR quantizers by adjusting the quantization range [28, 54] or modeling the feature distribution [16, 34] for activations, assigning a fixed bit for diverse image regions. However, these methods overlook that the accuracy degradation from quantization can vary for different contents, where some are more sensitive to quantization, thus showing a worse tolerance for low bits.
@@
To address this limitation, Hong et al. [15] propose content-aware dynamic quantization (CADyQ) which employs trainable bit selectors to measure the image and layer sensitivities for quantization simultaneously, as illustrated in Fig. 1(a). Nevertheless, incorporating such selectors into each layer will cause additional computational costs, particularly pronounced in deep networks. Several methods [39, 25] improve the trained selectors in CADyQ by exploring different image characteristics of patches, which conduct once more patch-wise quantization to tackle the image sensitivity. Though some advancements have been made, such a layer-wise bit-width adaption in response to varying patches can introduce disturbances to the inter-layer relations within original models to some extent, which leads to disparities in the representations, consequently compromising the reconstruction after quantization.
@@
These observations prompt us to consider a key question: Can we straightly adapt quantization with the awareness of image contents while avoiding layer sensitivity? In this context, deviating from existing methods, we rethink the quantization principle from two perspectives: 1) Granular characteristic, where fine-granularity representations reveal the texture complexity of local regions and coarse ones express structural semantics of the overall scene; 2) Entropy statistic, which reflects the average information density and the complexity of pixel distributions given patches [37], correlated with the image quality. Therefore, we propose a distinctive approach, dubbed Granular-DQ, which conducts low-bit dynamic quantization by harnessing the multi-granularity clues of diverse image contents to achieve efficient yet effective quantized SR models.
@@
Granular-DQ consists of two sequential policies: one to conduct granularity-aware bit allocation for all the patches and the other is fine-grained bit-width adaption based on the entropy (see Fig. 1(b)). For the former, we design a granularity-bit controller (GBC) that constructs a hierarchy of coarse-to-fine granularity representations for each patch. GBC then assigns an appropriate level of granularity to each patch, contingent upon its desired contribution percentage to the entire image, and aligns this with potential quantization bit-widths, enabling a tailored bit allocation. However, since Granular-DQ contains no bit constraint as CADyQ, relying solely on the GBC for quantization will force the network to be optimized toward reconstruction accuracy with pixel-wise supervision, leading to excessively high bits on some patches. To alleviate this, we present an entropy-based fine-tuning approach on the premise of GBC, making a fine-grained bit adjustment for the patches less quantized. We capture generalized distribution statistics of the entropy across large-scale data, providing approximate entropy thresholds to establish an entropy-to-bit (E2B) mechanism. The resultant entropy thresholds are then dynamically calibrated and fine-tuned by exploiting the entropy of calibration patches as the adaption factor, achieving a more precise bit assignment. Experiments on representative CNN- and transformer-based SR models demonstrate the superiority of Granular-DQ in the trade-off between accuracy and quantization efficiency over recent state-of-the-art methods. The main contributions are summarized as follows:
@@
2 Related Work
2.1 Single Image Super-Resolution.
Recent progress in CNNs has critically advanced the field of SISR, enhancing image quality and detail restoration significantly [9, 31]. However, the intensive computational demands of CNNs [9, 38, 51, 18, 26], transformer-based [30, 33, 6] and diffusion-based models [35, 36] limit their use in mobile and embedded systems. Efforts to mitigate computational complexity have spanned several dimensions, research has focused on several strategies, including lightweight architecture implementation [8, 43], knowledge distillation [19, 52], network pruning [53], re-parameterization [45], and parameter sharing [4]. Additionally, some adaptive networks have been investigated to refine both performance and efficiency dynamically [4, 44], highlighting the ongoing pursuit of an optimal balance between resource occupation and SR performance. However, apart from the computational complexity, the obstacle of memory storage imposed by floating-point operations also limits the usage of existing SR models. This work applies the network quantization technique for this purpose.
@@ 2.2 Network Quantization
Network quantization has emerged as an effective solution that transforms 32-bit floating point values into lower bits [55, 7, 56, 11, 3, 29] to improve the network efficiency, which can be divided into quantization-aware training (QAT) and post-training quantization (PTQ) methods. QAT [55, 7, 11, 3] integrates the quantization process into the training of networks, performing quantization adaption with complete datasets. PTQ methods [29, 46] often require a small calibration dataset to determine quantization parameters without retraining, which enables fast deployment on various devices. Recently, some methods introduce mixed-precision [10, 49] or dynamic quantization [32] into the above two paradigms, which allows for the automatic selection of the quantization precision of each layer. Though network quantization has been predominantly applied in various high-level tasks, its potential in SISR has not been fully exploited.
2.3 Quantization for Super-Resolution Networks
@@
Unlike high-level vision tasks, SISR presents unique challenges due to its high sensitivity to precision loss [28, 42, 16, 13]. PAMS [28] introduces the parameterized max scale scheme, which quantizes both weights and activations of the full-precision SR networks to fixed low-bit ones. DDTB [54] tackles the quantization of highly asymmetric activations by a layer-wise quantizer with dynamic upper and lower trainable bounds. DAQ [16] and QuantSR [34] study the influence of the parameter distribution in quantization, continuing to narrow the performance gap to full-precision networks. Recently, some attempts adopt dynamic quantization, which exploits the quantization sensitivity of layers and images, e.g. gradient magnitude [15], edge score [39], or cross-patch similarity [25], have demonstrated promising achievements. AdaBM [14] accelerates the adaptive quantization by separately processing image-wise and layer-wise bit-width adaption on the fly. In contrast, our method exploits the granularity and information density inherent in images to conduct dynamic quantization. It dispenses with the conventional need for layer sensitivity while being responsive to local contents, devising a distinctive patch-wise and layer-invariant dynamic quantization principle, which achieves superior performance and generalization ability for both CNN and transformer models.
3 Proposed Method
3.1 Preliminaries
@@
In most cases, converting the extensive floating-point calculations into operations that use fewer bits within CNNs involves quantizing the input features and weights at convolutional layers [23]. In the quantized SR network, given a quantizer $\mathcal{Q}$ in a symmetric mode, the function $\mathcal{Q}_{b}(\cdot)$ is applied to the input $\hat{x}_{k}$ of the $k$-th convolutional layer, transforming $x_{k}$ into its quantized counterpart $\hat{x}_{k}$ with a lower bit-width $b$, as expressed in the following formula
@@
where $\operatorname{clip}(\cdot)=max(min(x_{k},a),-a)$ confines $x_{k}$ within $[-a,a]$. $a$ denotes the maximum of the absolute value of $x$ [47] or derived from the moving average of max values across batches [42]. Additionally, $r_{b}$ serves as the mapping function that scales inputs of higher precision down to their lower bit equivalents, defined as $r_{b}=\frac{a}{2^{b-1}-1}$. Specially, the non-negative values after ReLU are truncated to $[0,a]$ and $r_{b}=\frac{a}{2^{b}-1}$. For weight quantization, given the $k$-th convolutional layer weight $w_{k}$, the quantized weight $\hat{w}_{i}$ can be formulated as follows
@@
Different from activations, the weights are quantized with fixed bit-width following [28, 15].
3.2 Motivation
Recent advances [40, 15, 39, 25, 14] have demonstrated the benefits of considering the quantization sensitivity of layers and image contents in SR quantization. Taking CADyQ [15] for example, it applies a trainable bit selector to determine the proper bit-width and quantization level for each layer and a given local image patch based on the feature gradient magnitude. In our analysis, we compute the average bit-width and the quantization error measured by MSE between the reconstructions of the quantized model (via CADyQ) and the original high-precision model (EDSR) on the Test2K dataset. Fig. 3(a) reveals that the majority of patches fall within a 6-bit to 8-bit range, accompanied by a relatively elevated MSE. Furthermore, we present t-SNE maps for various quantized layers and the final layer in Figs. 3(c)-(d). Firstly, it is evident that the distribution of different layers quantized by CADyQ is markedly more scattered than that of the original model, as depicted in Fig. 3(c). Secondly, on the final layer, the features from the CADyQ-quantized model exhibit a distinct vertical pattern, which is notably at odds with the structure of the original model’s feature points (Fig. 3(d)). In our investigation, CABM actually exhibited similar findings, although it fine-tunes the CADyQ model based on edge scores. These results indicate that: 1) Simply relying on image edge information is suboptimal for the trade-off between quantization efficiency and error; 2) The bit allocation for each layer in response to varying patches can introduce disturbances to the inter-layer relations within original models leading to disparities in the representations.
Figure 3: Analysis of the quantization efficiency, quantization error, and feature distribution in t-SNE on CADyQ and our Granular-DQ. (a) and (b) illustrate the quantization efficiency v.s. quantization error trade-off; (c) and (d) visualize the feature distribution of two resultant models and compare with the corresponding original one (Float32: EDSR).
Based on the above analysis, this work aims to design a dynamic quantization approach for diverse image contents while maintaining the representation ability of the original model. To this end, we rethink the image characteristics related to image quality from the granularity and information density. As we know, the fine-granularity representations reveal the texture complexity of local regions, while coarse ones express structural semantics of the overall scene. Besides, according to Shannon’s Second Theorem [37], the entropy statistic reflects the average information density and the complexity of pixel distributions given patches, which is directly correlated to the image quality. Therefore, we propose Granular-DQ, a markedly different method that fully explores the granularity and entropy statistic of images to quantization adaption. Granular-DQ contains two sequential steps: 1) granularity-aware bit allocation for all the patches and 2) entropy-based fine-grained bit-width adaption for the patches less quantized by 1). In this way, we can see that the bit-width allocation by Graular-DQ is sparser than CADyQ, where a majority of patches are lower than 5bit with only a few patches at high bit-width (Fig. 3 (b)). Moreover, the feature distribution of the layers quantized by our method is closer to that of the original model (Fig. 3(c)-(d)). These validate that our Granular-DQ enables low-bit and layer-invariant quantization.
3.3 Granular-DQ for SISR
The proposed Granular-DQ aims to cultivate a layer-invariant SR quantization approach that enables dynamic quantization of existing SR models for varying image contents with the awareness of multi-granularity clues. The overall pipeline is shown in Fig. 2, which contains two steps: 1) granularity-aware bit allocation by the granularity-bit controller (GBC) and 2) entropy-based fine-grained bit-width adaption on the patches allocated with high bits in GBC based on an entropy-to-bit (E2B) mechanism.
@@ 3.3 Granular-DQ for SISR
Granularity-Bit Controller. Given an image $X$, as shown in Fig. 4, the GBC first encodes it into hierarchical feature $\mathbf{Z}=\mathcal{E}(X)$ by the encoder $\mathcal{E}$, where $\mathbf{Z}={Z_{1},Z_{2},...,Z_{D}}$ via $D-1$ downsampling operations. Note that the resolution from $Z_{1}$ to $Z_{D}$ decreases progressively, where the largest $Z_{1}$ corresponds to the finest-granularity feature and the smallest $Z_{D}$ denotes the coarsest-granularity one (i.e. $D$ granularities), forming multi-granularity representations for $X$. We implement GBC with the Gumbel-Softmax, a differentiable sampling scheme [20], to adaptively measure the proportional contribution of all patches to the entire image, and align this with potential quantization bit-widths. To be specific, all the granularity features are group normalized and then average pooled to the coarsest granularity, i.e., with the same resolution of $Z_{D}$, denoted by $\mathbf{\hat{Z}}={\hat{Z}_{1},\hat{Z}_{2},...,\hat{Z}_{D}}$. We concatenate $\mathbf{\hat{Z}}$ along the channel dimension and squeeze the multi-granularity information by global average pooling $GAP(\cdot)$ to generate a channel-wise statistics $\mathbf{S}$ of $X$, formulated by
@@
Inspired by the end-to-end discrete methodology in [48], the fixed decision typically dictated by Eq.(5) is substituted with a probabilistic sampling approach. It hinges on the utilization of a categorical distribution characterized by unnormalized log probabilities, from which discrete gating indices are derived by integrating a noise sample $\sigma_{n}$, originating from the standard Gumbel distribution $\mathrm{Gumbel}(0,1)$:

### v2

@@ Abstract
Dynamic quantization has attracted rising attention in image super-resolution (SR) as it expands the potential of heavy SR models onto mobile devices while preserving competitive performance. Existing methods explore layer-to-bit configuration upon varying local regions, adaptively allocating the bit to each layer and patch. Despite the benefits, they still fall short in the trade-off of SR accuracy and quantization efficiency. Apart from this, adapting the quantization level for each layer individually can disturb the original inter-layer relationships, thus diminishing the representation capability of quantized models. In this work, we propose Granular-DQ, which capitalizes on the intrinsic characteristics of images while dispensing with the previous consideration for layer sensitivity in quantization. Granular-DQ conducts a multi-granularity analysis of local patches with further exploration of their information densities, achieving a distinctive patch-wise and layer-invariant dynamic quantization paradigm. Specifically, Granular-DQ initiates by developing a granularity-bit controller (GBC) to apprehend the coarse-to-fine granular representations of different patches, matching their proportional contribution to the entire image to determine the proper bit-width allocation. On this premise, we investigate the relation between bit-width and information density, devising an entropy-to-bit (E2B) mechanism that enables further fine-grained dynamic bit adaption of high-bit patches. Extensive experiments validate the superiority and generalization ability of Granular-DQ over recent state-of-the-art methods on various SR models. Code and supplementary statement can be found at https://github.com/MmmingS/Granular-DQ.git.
@@
Figure 1: Visual comparison of (a) previous dynamic quantization pipeline (Hong et al. 2022a) that adapt the bit allocation for layers and patches simultaneously and (b) our Granular-DQ pipeline conducts patch-wise and layer-invariant dynamic quantization, which contains two steps: 1) granularity-aware bit allocation and 2) fine-grained bit-width adaption based on the entropy statistics. Our method recovers a better SR image with a lower average bit.
Introduction
@@
Single image super-resolution (SISR) has been a fundamental task in the computer vision community, aiming to recover high-resolution (HR) images from corrupted low-resolution (LR) input. Recently, from the pioneering deep learning-based method (Dong et al. 2014), convolutional neural networks (CNN) (Dong et al. 2014; Kim, Lee, and Lee 2016; Shi et al. 2016; Zhang et al. 2018; Ahn, Kang, and Sohn 2018; Li et al. 2022) and transformers (Liang et al. 2021; Lu et al. 2022; Zhang et al. 2022; Chen et al. 2023) have dominated SISR. While the SR performance continues to achieve breakthroughs, the model complexity of later methods also increases constantly, which limits their practical applications, especially tackling large-size images (e.g. 2K and 4K). This raises interest in compressing deep SR models to unlock their potential on resource-constrained devices.
@@
Model quantization (Zhou et al. 2016) has emerged as a promising technology that reduces both computational overhead and memory cost with minimal performance sacrifice, where the effectiveness has been demonstrated in a wide range of high-level tasks (Zhou et al. 2016; Choi et al. 2018; Bhalgat et al. 2020; Chen et al. 2021; Gao et al. 2022; Luo et al. 2023). Some prior works design SR quantizers by adjusting the quantization range (Li et al. 2020; Zhong et al. 2022) or modeling the feature distribution  (Hong et al. 2022b; Qin et al. 2024) for activations, assigning a fixed bit for diverse image regions. However, these methods overlook that the accuracy degradation from quantization can vary for different contents, where some are more sensitive to quantization, thus showing a worse tolerance for low bits.
@@
To address this limitation, Hong et al. (Hong et al. 2022a) propose content-aware dynamic quantization (CADyQ) which employs trainable bit selectors to measure the image and layer sensitivities for quantization simultaneously, as illustrated in Figure 1(a). Nevertheless, incorporating such selectors into each layer will cause additional computational costs, particularly pronounced in deep networks. Several methods (Tian et al. 2023; Lee, Yoo, and Jung 2024) improve the trained selectors in CADyQ by exploring different image characteristics of patches, which conduct once more patch-wise quantization to tackle the image sensitivity. Though some advancements have been made, such a layer-wise bit-width adaption in response to varying patches can introduce disturbances to the inter-layer relations within original models to some extent, which leads to disparities in the representations, consequently compromising the reconstruction after quantization.
@@
These observations prompt us to consider a key question: Can we straightly adapt quantization with the awareness of image contents while avoiding layer sensitivity? In this context, deviating from existing methods, we rethink the quantization principle from two perspectives: 1) Granular characteristic, where fine-granularity representations reveal the texture complexity of local regions and coarse ones express structural semantics of the overall scene; 2) Entropy statistic, which reflects the average information density and the complexity of pixel distributions given patches (Shannon 1948), correlated with the image quality. Therefore, we propose a distinctive approach, dubbed Granular-DQ, which conducts low-bit dynamic quantization by harnessing the multi-granularity clues of diverse image contents to achieve efficient yet effective quantized SR models.
@@
Granular-DQ consists of two sequential policies: one to conduct granularity-aware bit allocation for all the patches and the other is fine-grained bit-width adaption based on the entropy (see Figure 1(b)). For the former, we design a granularity-bit controller (GBC) that constructs a hierarchy of coarse-to-fine granularity representations for each patch. GBC then assigns an appropriate level of granularity to each patch, contingent upon its desired contribution percentage to the entire image, and aligns this with potential quantization bit-widths, enabling a tailored bit allocation. However, since Granular-DQ contains no bit constraint as CADyQ, relying solely on the GBC for quantization will force the network to be optimized toward reconstruction accuracy with pixel-wise supervision, leading to excessively high bits on some patches. To alleviate this, we present an entropy-based fine-tuning approach on the premise of GBC, making a fine-grained bit adjustment for the patches less quantized. We capture generalized distribution statistics of the entropy across large-scale data, providing approximate entropy thresholds to establish an entropy-to-bit (E2B) mechanism. The resultant entropy thresholds are then dynamically calibrated and fine-tuned by exploiting the entropy of calibration patches as the adaption factor, achieving a more precise bit assignment. Experiments on representative CNN- and transformer-based SR models demonstrate the superiority of Granular-DQ in the trade-off between accuracy and quantization efficiency over recent state-of-the-art methods. The main contributions are summarized as follows:
@@
Related Work
Single Image Super-Resolution.
Recent progress in CNNs has critically advanced the field of SISR, enhancing image quality and detail restoration significantly  (Dong et al. 2014; Lim et al. 2017). However, the intensive computational demands of CNNs (Dong et al. 2014; Shi et al. 2016; Zhang et al. 2018; Hui, Wang, and Gao 2018; Li, Bai, and Zhao 2020), transformer-based (Liang et al. 2021; Lu et al. 2022; Chen et al. 2023) and diffusion-based models (Rombach et al. 2022; Saharia et al. 2023) limit their use in mobile and embedded systems. Efforts to mitigate computational complexity have spanned several dimensions, research has focused on several strategies, including lightweight architecture implementation (Chu et al. 2021; Wang et al. 2021b), knowledge distillation (Hui et al. 2019; Zhang et al. 2021a), network pruning (Zhang et al. 2021b), re-parameterization (Wang, Dong, and Shan 2022), and parameter sharing (Chen et al. 2022). Additionally, some adaptive networks have been investigated to refine both performance and efficiency dynamically (Chen et al. 2022; Wang et al. 2022), highlighting the ongoing pursuit of an optimal balance between resource occupation and SR performance. However, apart from the computational complexity, the obstacle of memory storage imposed by floating-point operations also limits the usage of existing SR models. This work applies the network quantization technique for this purpose.
@@
Network quantization has emerged as an effective solution that transforms 32-bit floating point values into lower bits (Zhou et al. 2016; Choi et al. 2018; Zhuang et al. 2018; Esser et al. 2019; Bhalgat et al. 2020; Li et al. 2021) to improve the network efficiency, which can be divided into quantization-aware training (QAT) and post-training quantization (PTQ) methods. QAT (Zhou et al. 2016; Choi et al. 2018; Esser et al. 2019; Bhalgat et al. 2020) integrates the quantization process into the training of networks, performing quantization adaption with complete datasets. PTQ methods (Li et al. 2021; Wei et al. 2022) often require a small calibration dataset to determine quantization parameters without retraining, which enables fast deployment on various devices. Recently, some methods introduce mixed-precision (2019) or dynamic quantization (2022) into the above two paradigms, which allows for the automatic selection of the quantization precision of each layer. Though network quantization has been predominantly applied in various high-level tasks, its potential in SISR has not been fully exploited.
Quantization for Super-Resolution Networks
@@
Unlike high-level vision tasks, SISR presents unique challenges due to its high sensitivity to precision loss (Li et al. 2020; Wang et al. 2021a; Hong et al. 2022b; Hong and Lee 2023). PAMS (Li et al. 2020) introduces the parameterized max scale scheme, which quantizes both weights and activations of the full-precision SR networks to fixed low-bit ones. DDTB (Zhong et al. 2022) tackles the quantization of highly asymmetric activations by a layer-wise quantizer with dynamic upper and lower trainable bounds. DAQ (Hong et al. 2022b) and QuantSR (Qin et al. 2024) study the influence of the parameter distribution in quantization, continuing to narrow the performance gap to full-precision networks. Recently, some attempts adopt dynamic quantization, which exploits the quantization sensitivity of layers and images, e.g. gradient magnitude (Hong et al. 2022a), edge score (Tian et al. 2023), or cross-patch similarity (Lee, Yoo, and Jung 2024), have demonstrated promising achievements. AdaBM (Hong and Lee 2024) accelerates the adaptive quantization by separately processing image-wise and layer-wise bit-width adaption on the fly. In contrast, our method exploits the granularity and information density inherent in images to conduct dynamic quantization. It dispenses with the conventional need for layer sensitivity while being responsive to local contents, devising a distinctive patch-wise and layer-invariant dynamic quantization principle, which achieves superior performance and generalization ability for both CNN and transformer models.
Proposed Method
Preliminaries
@@
In most cases, converting the extensive floating-point calculations into operations that use fewer bits within CNNs involves quantizing the input features and weights at convolutional layers (Krishnamoorthi 2018). In the quantized SR network, given a quantizer $\mathcal{Q}$ in a symmetric mode, the function $\mathcal{Q}_{b}(\cdot)$ is applied to the input $\hat{x}_{k}$ of the $k$-th convolutional layer, transforming $x_{k}$ into its quantized counterpart $\hat{x}_{k}$ with a lower bit-width $b$, as expressed in the following formula
@@
where $\operatorname{clip}(\cdot)=max(min(x_{k},a),-a)$ confines $x_{k}$ within $[-a,a]$. $a$ denotes the maximum of the absolute value of $x$ (Wu et al. 2020) or derived from the moving average of max values across batches (Wang et al. 2021a). Additionally, $r_{b}$ serves as the mapping function that scales inputs of higher precision down to their lower bit equivalents, defined as $r_{b}=\frac{a}{2^{b-1}-1}$. Specially, the non-negative values after ReLU are truncated to $[0,a]$ and $r_{b}=\frac{a}{2^{b}-1}$. For weight quantization, given the $k$-th convolutional layer weight $w_{k}$, the quantized weight $\hat{w}_{i}$ can be formulated as follows
@@
Different from activations, the weights are quantized with fixed bit-width following (Li et al. 2020; Hong et al. 2022a).
Granular-DQ for SISR
The proposed Granular-DQ aims to cultivate a layer-invariant SR quantization approach that enables dynamic quantization of existing SR models for varying image contents with the awareness of multi-granularity clues. The overall pipeline is shown in Figure 2, which contains two steps: 1) granularity-aware bit allocation by the granularity-bit controller (GBC) and 2) entropy-based fine-grained bit-width adaption on the patches allocated with high bits in GBC based on an entropy-to-bit (E2B) mechanism.
@@
Granularity-Bit Controller. Given an image $X$, as shown in Figure 3, the GBC first encodes it into hierarchical feature $\mathbf{Z}=\mathcal{E}(X)$ by the encoder $\mathcal{E}$, where $\mathbf{Z}={Z_{1},Z_{2},...,Z_{D}}$ via $D-1$ downsampling operations. Note that the resolution from $Z_{1}$ to $Z_{D}$ decreases progressively, where the largest $Z_{1}$ corresponds to the finest-granularity feature and the smallest $Z_{D}$ denotes the coarsest-granularity one (i.e. $D$ granularities), forming multi-granularity representations for $X$. We implement GBC with the Gumbel-Softmax, a differentiable sampling scheme (Jang, Gu, and Poole 2017), to adaptively measure the proportional contribution of all patches to the entire image, and align this with potential quantization bit-widths. To be specific, all the granularity features are group normalized and then average pooled to the coarsest granularity, i.e., with the same resolution of $Z_{D}$, denoted by $\mathbf{\hat{Z}}={\hat{Z}_{1},\hat{Z}_{2},...,\hat{Z}_{D}}$. We concatenate $\mathbf{\hat{Z}}$ along the channel dimension and squeeze the multi-granularity information by global average pooling $GAP(\cdot)$ to generate a channel-wise statistics $\mathbf{S}$ of $X$, formulated by
@@
Inspired by the end-to-end discrete methodology in (Xie et al. 2020), the fixed decision typically dictated by Eq.(5) is substituted with a probabilistic sampling approach. It hinges on the utilization of a categorical distribution characterized by unnormalized log probabilities, from which discrete gating indices are derived by integrating a noise sample $\sigma_{n}$, originating from the standard Gumbel distribution $\mathrm{Gumbel}(0,1)$: